# How (Mis)calibrated is Your Federated CLIP and What To Do About It?

### v1

@@ 4.4 What makes $\text{FL}^{2}\text{oRA}$ effective?
In this section, we provide a high-level, largely qualitative discussion of why $\text{FL}^{2}\text{oRA}$ may be particularly effective for federated CLIP fine-tuning. We emphasize that this is an interpretive account grounded in prior work and our empirical trends, rather than a mechanism that we have explicitly validated through targeted ablations.
During local training, each client updates only the low-rank LoRA parameters while the original CLIP weights and classifier remain frozen; the server then aggregates these adapter parameters using FedAvg. As illustrated in Fig. 1 and Tab. 3, 4, and 5, this lightweight modification consistently improves both accuracy and calibration relative to existing CLIP-based FL methods across all benchmarks.
Implicit ensembling via FedAvg. FedAvg can be viewed as a form of implicit ensembling or weight-space bagging: each client trains on its own data partition, and the server averages their parameters. Ensembling is well known to improve calibration [lakshminarayanan2017simple, Ashukha2020Pitfalls], but its benefits rely on the individual models having reasonably compatible confidence scales i.e., none of them being extremely overconfident or poorly calibrated [yong2024spurious]. From this perspective, earlier FL methods that fully fine-tune all weights may produce client models whose logits are very differently scaled, limiting the extent to which FedAvg can act as a stabilizing ensemble [nucfl, peng2024fedcal]. In contrast, if each client’s local model stays closer to a well-calibrated pre-trained solution, averaging their parameters is more likely to yield a calibrated global model [yong2024spurious].
LoRA helps preserve calibration. LoRA typically leaves most pre-trained weights unchanged and restricts adaptation to a low-rank update [shi2024long]. Prior work suggests that such updates can remain closer to the pre-trained solution in weight space and preserve its behavior more faithfully than standard regularization alone [biderman2024lora]. From a calibration standpoint, this may help avoid large shifts in weight norms and logit scales, which are known contributors to miscalibration [shi2024long, minderer2021revisiting]. By constraining local adaptation to a low-rank subspace, $\text{FL}^{2}\text{oRA}$ encourages each client to stay within a neighborhood where CLIP is already reasonably calibrated, while still allowing meaningful task-specific and client-specific adaptation.
Putting these observations together, we hypothesize that $\text{FL}^{2}\text{oRA}$ benefits from a favorable combination of (i) CLIP’s strong zero-shot calibration as a starting point, (ii) the stability and norm control induced by LoRA updates, and (iii) the ensemble-like regularization effect of FedAvg over heterogeneous clients. While our experiments do not isolate these factors individually, this perspective is consistent with the empirical finding that $\text{FL}^{2}\text{oRA}$ not only preserves but often improves calibration while simultaneously boosting accuracy in federated learning.
@@ 4.5 Ablation Studies
We perform a thorough ablation analysis on the in-distribution setting using the CIFAR-100 dataset. For each ablation study we report the accuracy and ECE metrics.
Sensitivity to data distribution. We study the impact of varying $\alpha$, the concentration parameter of the Dirichlet distribution, which controls the overlap of classes among clients. A low value of $\alpha$ signifies a low overlap among clients (ie, non-IID), and a high value of alpha signifies a high overlap among clients (ie, IID). From Fig. 2 we observe that for our proposed $\text{FL}^{2}\text{oRA}$, both accuracy and ECE remain fairly stable at varying levels of heterogeneity. Although competitor methods, such as FedPGP and FedPHA, exhibit improved ECE with respect to the simple baseline ZS-CLIP, they fluctuate with different values of $\alpha$.
@@
Figure 2: Effect of varying $\alpha$ of Dirichlet distribution. We report (a) Accuracy and (b) ECE metrics on CIFAR-100 dataset.
(a) Accuracy
(b) ECE
Figure 3: Effect of varying communication rounds. We report (a) Accuracy and (b) ECE metrics on CIFAR-100 dataset.
@@
(a) Accuracy
(b) ECE
Figure 4: Ablation of rank of LoRA metrices as (a) Accuracy and (b) ECE metrics on CIFAR-100 dataset. To be noted, ‘T’ and ‘V’ refer to text and vision respectively.
Sensitivity to rank in $\text{FL}^{2}\text{oRA}$. We study the impact of rank $r$ of LoRA, as it controls how much capacity the adapter has to modify the base CLIP model. A higher rank means more expressive updates, at the cost of increasing the computation and communication overhead, and vice versa. In Fig. 4 we report the different variants of $\text{FL}^{2}\text{oRA}$, where we observe that the performance of $\text{FL}^{2}\text{oRA}$ remains stable in varying ranks. We postulate that performance remains stable because even a lower rank is sufficient to adapt to the new knowledge contained in CIFAR-100. However, for a more complex dataset, a higher rank might be necessary.
Sensitivity to PEFT strategies in $\text{FL}^{2}\text{oRA}$. In Fig. 5 we investigate the impact of different PEFT methods including LayerNorm [layernorm] and BitFit [bitfit], alongside LoRA [lora] in $\text{FL}^{2}\text{oRA}$. Across all experiments, jointly fine-tuning both the text and vision CLIP encoders consistently outperforms unimodal tuning (text or vision) in terms of both accuracy and ECE. Notably, $\text{FL}^{2}\text{oRA}$ consistently surpasses the LayerNorm and BitFit under all modality tuning configurations, highlighting its effectiveness in maintaining high accuracy while preserving calibration in a FL setup.
(a) Accuracy
(b) ECE
Figure 5: Ablation of PEFT strategies. as (a) Accuracy and (b) ECE metrics on CIFAR-100 dataset.
@@ 5 Conclusions
In this work, we investigated the an unexplored question of how fine-tuning CLIP in a FL setting affects model calibration. Our study reveals that TPT-based methods that perform reasonably in offline settings exhibit worse calibration when operating in FL. While standard in-training calibration techniques offer limited improvements, our analysis suggests that the primary driver of miscalibration is the way prompts are adapted under client heterogeneity. Encouragingly, we find that PEFT approaches, such as LoRA can help. Therefore, we present $\text{FL}^{2}\text{oRA}$, and show that it naturally yield to well-calibrated models in FL, substantially reducing the need for additional calibration procedures. By exploring the link between parameter-efficient fine-tuning and FL, our work opens new directions for addressing calibration in CLIP within the federated setting.
Table of Supplementary Contents
Appendix A Experimental setup
A.1 Benchmarks
@@
For the base-to-new generalization and domain generalization settings, we consider more challenging datasets that assess the model’s ability to transfer knowledge to unseen classes or in between multiple domains. In the base-to-new setting, we choose Caltech101, Flowers102, OxfordPets, Food101, and DTD datasets in few-shot training set up. In the domain generalization (DG) setting, we use traditional DG benchmarks having four domains to each i.e. PACS (art painting, cartoon, photo and sketch), OfficeHome (art, clipart, product and realworld), and VLCS (caltech, labelme, pascal-voc and sun), providing distribution shifts.
@@
Figure 6: Federated data partitioning under different Dirichlet concentration parameter $\alpha$. In each of the subplots, each row represents a client and each cell corresponds to a data sample whose color indicates its class label. The sub figures are represented as, (a) Sort and partition setting with two fixed classes per client, (b) to (f) Non-IID splits with Dirichlet distribution, having $\alpha\!\to\!\infty$, $\alpha=100,1,0.5$, and $\alpha\!\to\!0$. Larger $\alpha$ produces nearly IID client distributions, whereas smaller $\alpha$ concentrates samples from few classes only within each client with increasingly skewed and heterogenous partitions.
A.2 Federated learning settings
We describe in detail the three personalized FL settings i.e. in-distribution, domain generalization and base-to-new generalization tackled in this work.
[-] In-distribution & Domain generalization. For both settings, we distribute the classes across clients in an overlapping non-IID manner, as similar to [fedotp]. To simulate controllable non-IID data heterogeneity across clients, we adopt the standard Dirichlet-based partitioning strategy. Given a dataset with $C$ classes and $N$ clients, each client’s class proportion vector, $\mathbf{q}_{n}=(q_{n1},\dots,q_{nK})$ is sampled from a symmetric Dirichlet distribution [bayesian] parameterized by concentration $\alpha$:
@@ A.2 Federated learning settings
where $|\mathcal{D}_{n}|$ denotes the local dataset size. The concentration parameter $\alpha$ directly controls the heterogeneity of the federated setting i.e. large $\alpha$ values (e.g., $\alpha\!\rightarrow\!\infty$) yield nearly IID client distributions, while small $\alpha$ values (e.g., $\alpha\!\rightarrow\!0$) produce extremely skewed, label-concentrated clients. Figure 6 visualizes this progression, showing how the distribution transitions from highly mixed to strongly partitioned as $\alpha$ decreases. We keep $\alpha=0.5$ for both the federated settings to maintain the overlapping non-IID in a balanced manner. For in-distribution, we choose 100 clients over 10% participation rate, whereas two clients per domain with 100% participation rate in domain generalization setting.
@@
A.3 Evaluation protocol and metrics
Evaluation protocol. We evaluate the methods in three aforementioned FL settings in a personalized FL setup [fedotp, fedpha, fedpgp], where we compute the metrics on the test set in each client and then take an average over all the clients. To be noted that client participation in training may vary according to the participation rate, but inference is performed on all clients regardless of their participation.
Metrics. In our FL experiments, we use the accuracy metric for model performance and ECE, MCE and ACE metrics for model calibration.
@@ A.3 Evaluation protocol and metrics
[-] Expected Calibration Error (ECE). ECE [ece] measures the average discrepancy between a model’s predicted confidence and its actual accuracy. The prediction confidences are partitioned into $G$ bins, and for each bin $b_{g}$, the absolute difference between the empirical accuracy $\text{acc}(b_{g})$ and the mean predicted confidence $\text{conf}(b_{g})$ is computed. These bin-wise errors are then weighted by the proportion of samples in each bin, yielding a global estimate of how well the model’s confidence reflects the true likelihood of correctness. Lower ECE values indicate better-calibrated predictions. The ECE metric can be defined as:
@@
[-] Maximum Calibration Error (MCE). MCE [ece] quantifies the worst-case miscalibration across all bins. Rather than averaging, MCE takes the maximum absolute gap between accuracy and confidence over the $G$ bins. This highlights extreme local over- or under-confidence that may be masked by aggregate metrics such as ECE. A lower MCE value indicates that the model avoids severe calibration failures in any confidence region. The ECE metric can be written as:
@@
[-] Average Calibration Error (ACE). ACE [ace] assesses calibration by averaging the absolute accuracy-confidence discrepancies uniformly across all $G$ bins. Unlike ECE, which weights bins by the number of samples they contain, ACE treats each bin equally, making it more sensitive to miscalibration in sparsely populated regions of the confidence distribution. Lower ACE values reflect more consistent calibration performance across different confidence levels. The ECE metric can be defined as:
$\text{ACE}=\frac{1}{G}\sum^{G}_{g=1}\mid\text{acc}(b_{g})-\text{conf}(b_{g})\mid,$ (7)
A.4 Calibration baselines
In our experimental settings, we consider two standalone calibration methods DCA [dca] and MDCA [mdca] as additional plug-ins in the federated methods.
[-] Difference-between-Confidence-Accuracy (DCA). DCA [dca] provides a minibatch-level calibration objective by penalizing the discrepancy between the model’s average predicted confidence score corresponding to the true class and its average empirical accuracy for all the samples in a minibatch. For each prediction corresponding to the $i$-th sample, the method computes the correctness indicator $c_{i}$ (which equals $1$ if the prediction is correct and 0 otherwise) and the predicted confidence of the true label $s_{i}=p(\hat{y}_{i})$. By minimizing this deviation, DCA encourages the model’s overall confidence level to align with its average true accuracy, thereby improving global calibration behavior. For a minibatch of size $m$, the loss is computed as:

### v2

@@ 4.4 What makes $\text{FL}^{2}\text{oRA}$ effective?
In this section, we provide a high-level, largely qualitative discussion of why $\text{FL}^{2}\text{oRA}$ may be particularly effective for federated CLIP fine-tuning. In the FL strategies, the final weight is a weighted average of the different learning rounds as written in Eq.˜2. Based on [izmailov2018averaging], at communication round $t$, we can decompose the output of the local model of client $n$ via a first-order Taylor expansion around the global model $\theta_{t}$:
$f_{\theta^{(n)}_{t}}(\mathbf{x}^{(i)})=f_{\theta_{t}}(\mathbf{x}^{(i)})+\bigl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\Delta_{n}\bigr\rangle+O\bigl(\|\Delta_{n}\|^{2}\bigr),$ (3)
where $\Delta_{n}=\theta^{(n)}_{t}-\theta_{t}$ is the local deviation from the global model, and $\nabla f_{\theta_{t}}(\mathbf{x}^{(i)})$ is the Jacobian of $f$ with respect to the parameters $\theta$, evaluated at $\theta_{t}$ and input $\mathbf{x}^{(i)}$.
By multiplying Eq.˜3 by the client weight $\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}$ and summing over all clients $n$, we obtain:
$\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\,f_{\theta^{(n)}_{t}}(\mathbf{x}^{(i)})=\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\Bigl[f_{\theta_{t}}(\mathbf{x}^{(i)})+\bigl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\Delta_{n}\bigr\rangle+O\bigl(\|\Delta_{n}\|^{2}\bigr)\Bigr].$ (4)
The left-hand side is precisely the weighted ensemble of the local model predictions, equivalent to Deep Ensemble [lakshminarayanan2017simple, yong2024spurious]. On the right-hand side, the first term simplifies as follows:
$\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\,f_{\theta_{t}}(\mathbf{x}^{(i)})=f_{\theta_{t}}(\mathbf{x}^{(i)}),$
since $\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}=1$ by definition. The linear term vanishes because:
$\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\bigl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\Delta_{n}\bigr\rangle=\biggl\langle\nabla f_{\theta_{t}}(\mathbf{x}^{(i)}),\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\Delta_{n}\biggr\rangle=0,$
where we used the fact that $\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\Delta_{n}=\mathbf{0}$. The remainder term is therefore
$O\biggl(\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\|\Delta_{n}\|^{2}\biggr),$ which is second-order small when the local deviations $\Delta_{n}$ remain small.
Hence, the Federated Learning global model prediction $f_{\theta_{t}}(\mathbf{x}^{(i)})$ is approximately equal to the weighted Deep Ensemble: $\sum_{n}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\,f_{\theta^{(n)}_{t}}(\mathbf{x}^{(i)}).$
Recent work has shown that Deep Ensembles yield superior calibration compared to single models [lakshminarayanan2017simple, ovadia2019can]. In addition, in our setting, we replace full fine-tuning by LoRA [hu2022lora], so the global model corresponds to a LoRA Ensemble [wang2023lora, muhlematter2024lora]. Finally, some works have shown that LoRA Ensemble provides a computationally efficient approximation to Deep Ensemble [wenbatchensemble, muhlematter2024lora]. Based on these results, we can therefore interpret $\text{FL}^{2}\text{oRA}$ as an effective approximation to a Deep Ensemble, inheriting many of its generalization and uncertainty quantification benefits while remaining finetuning efficient.
To understand what makes $\text{FL}^{2}\text{oRA}$ different from TPT-based FL methods, where there is also an ensembling operation due to FedAvg, we take a closer at CLIP classification mechanism. In CLIP inference, the visual features are obtained from the image encoder, and the “classifier” weights are constructed from the text embeddings. In TPT-based FL methods, the ensembling operation is effectively performed only on the so-called classifier weights, and therefore can be considered equivalent to a last-layer ensemble [legate2023guiding], where only the last layer is fine-tuned and ensembled. However, without explicit mechanism to encourage diversity, such as repulsion losses, last-layer ensemble often exhibit limited functional diversity [steger2024function] compared to Deep Ensembles. In contrast, techniques such as LoRA Ensemble [muhlematter2024lora, wang2023lora], Batch Ensembles [wenbatchensemble] or Packed-Ensembles [laurent2023packed], which often operate on the level of the feature extraction backbone, have a greater member diversity.
@@ 4.5 Ablation Studies
In this section, we perform a comprehensive ablation study on the in-distribution setting using the CIFAR-10/100 dataset.
Figure 2: Ablation on the LoRA rank and placement. (left) Accuracy and (right) ECE on CIFAR-100 for different LoRA ranks $r$ and encoder placements. ‘T’ and ‘V’ denote applying LoRA to the text and vision encoders, respectively. $\text{FL}^{2}\text{oRA}$ remains stable across different ranks and consistently outperforms the best TPT baseline, even at very low ranks.
$\text{FL}^{2}\text{oRA}$ learning capacity. One may attribute the superior performance of $\text{FL}^{2}\text{oRA}$ to the larger number of trainable parameters during fine-tuning. To examine this, we analyze the effect of the LoRA rank $r$, which controls the adapter capacity to modify the base CLIP model. We also ablate the location where LoRA is applied (text and/or image encoders), which changes the total number of trainable parameters. A higher rank enables more expressive updates but increases computation and communication overhead. As shown in Fig. 2, $\text{FL}^{2}\text{oRA}$ remains stable across different ranks and consistently outperforms the best TPT baseline; even with a rank as small as 2, $\text{FL}^{2}\text{oRA}$ still achieves better performance than TPT methods.
Table 7: Effect of PEFT variants on CIFAR-10 dataset. Best number is in bold, and second best is underlined.
Method Text-encoder Image-encoder Both encoders
#Tr. Params Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ #Tr. Params Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ #Tr. Params Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
CLIP-FT [radford2019language] 38.1M 92.31 5.16 1.66 4.34 86.2M 95.57 3.87 1.41 3.33 124.3M 95.76 3.44 1.25 2.48
AdaptFormer [adaptformer] 30.8K 92.51 5.39 1.87 3.92 46.1K 96.64 3.06 1.20 2.09 76.9K 96.80 3.00 1.16 1.93
LayerNorm [layernorm] 49.1K 90.42 5.56 1.76 4.70 75.3K 96.04 3.45 1.26 2.39 124.4K 96.03 3.49 1.30 2.21
BitFit [bitfit] 24.6K 92.96 4.67 1.53 3.66 39.9K 96.09 3.08 1.19 2.25 64.5K 96.49 3.14 1.10 2.18
DoRA [dora] 92.2K 92.65 5.08 1.64 4.06 138.2K 97.23 2.99 1.12 1.90 230.4K 97.19 3.05 1.08 1.85
AdaLoRA [adalora] 73.8K 88.37 8.61 2.33 7.92 110.7K 90.19 8.26 2.31 7.71 184.5K 89.04 8.47 2.33 8.02
VeRA [vera] 36.9K 88.50 8.54 2.34 7.91 55.3K 88.77 8.55 2.37 8.19 92.2K 88.93 8.30 2.25 7.82
$\text{FL}^{2}\text{oRA}$ (Ours) 73.7K 92.64 4.99 1.68 3.77 110.6K 97.08 2.94 1.14 1.98 184.3K 97.04 2.99 1.12 2.06
Superiority of $\text{FL}^{2}\text{oRA}$ over PEFT variants. In Table 7 we show that $\text{FL}^{2}\text{oRA}$ consistently provides a stronger balance between accuracy and calibration compared to existing PEFT approaches across different encoder tuning settings. Unlike several baselines that either improve accuracy at the expense of calibration or require more trainable parameters, $\text{FL}^{2}\text{oRA}$ maintains competitive predictive performance while producing more reliable confidence estimates. This trend is particularly evident when adapting the image encoder, where $\text{FL}^{2}\text{oRA}$ achieves the most favorable calibration behavior while remaining highly accurate. Even when both encoders are updated, the method preserves this balance without introducing significant parameter overhead. These results suggest that the proposed factorized low-rank formulation improves the effectiveness of parameter-efficient adaptation by capturing task-relevant updates while avoiding the over-confident predictions commonly observed in other PEFT strategies.
@@
Figure 3: Effect of varying communication rounds. We report (a) Accuracy and (b) ECE on CIFAR-100 dataset.
@@
Sensitivity to data distribution. We study the impact of varying $\beta$, the concentration parameter of the Dirichlet distribution, which controls the overlap of classes among clients. A low value of $\beta$ signifies a low overlap among clients (i.e., non-IID), and a high value indicates a high overlap among clients (i.e., IID). From Fig. 4 we observe that for our proposed $\text{FL}^{2}\text{oRA}$ both accuracy and ECE remain fairly stable at varying levels of heterogeneity. Although competitor methods, such as FedPGP and FedPHA, exhibit improved ECE with respect to the simple baseline ZS-CLIP, they fluctuate with different values of $\beta$.
Figure 4: Effect of varying $\beta$ of Dirichlet distribution. We report (left) Accuracy and (right) ECE on CIFAR-100 dataset.
Effect of post-hoc calibration technique. Temperature Scaling [guo2017calibration] is a widely used post-hoc calibration technique that rescales the network output (logits) by dividing them by a learned scalar parameter $\tau$, typically estimated on a validation set. Figure 5 illustrates how $\tau$ affects model calibration (ECE) on the CIFAR-10 and CIFAR-100 datasets [guo2017calibration]. The results showcase moderate temperatures yield better calibration, while too small or too large $\tau$ harms it. ZS-CLIP is highly sensitive to $\tau$ as ECE improves markedly near $\tau=0.5$ but degrades rapidly as $\tau$ increases, indicating that mild softening corrects overconfidence whereas excessive softening induces underconfident predictions. Among federated baselines, FedOTP and FedPGP tend to worsen as $\tau$ grows, suggesting a monotonic loss of calibration with over-smoothing, whereas PromptFL and FedPHA are comparatively stable with smaller ECE variation. However, our proposed $\text{FL}^{2}\text{oRA}$ maintains both the lowest and most stable ECE over a wide range of $\tau$, pointing to robustness to the temperature choice and reduced reliance on per-dataset tuning.
(a) CIFAR-10
(b) CIFAR-100
@@ 5 Conclusions
In this work, we investigated the an unexplored question of how fine-tuning CLIP in a FL setting affects model calibration. Our study reveals that TPT-based methods that perform reasonably in offline settings exhibit worse calibration when operating in FL. While standard in-training calibration techniques offer limited improvements, our analysis suggests that the primary driver of miscalibration is the we fine-tune CLIP. Therefore, we present $\text{FL}^{2}\text{oRA}$, and show that it naturally yield to well-calibrated models in FL, substantially reducing the need for additional calibration procedures. By exploring the link between PEFT and FL, our work opens new directions for addressing calibration in CLIP within the federated setting.
Appendix: How (Mis)calibrated is Your Federated CLIP and What To Do About It?
Appendix 0.A Experimental setup
0.A.1 Benchmarks
@@
For the base-to-new generalization and domain generalization settings, we consider more challenging datasets that assess the model’s ability to transfer knowledge to unseen classes or between multiple domains. In the base-to-new setting, we choose Caltech101, Flowers102, OxfordPets, Food101, and DTD datasets in a few-shot training setup. In the domain generalization (DG) setting, we use traditional DG benchmarks having four domains each i.e. PACS (art painting, cartoon, photo and sketch), OfficeHome (art, clipart, product and realworld), and VLCS (caltech, labelme, pascal-voc and sun), providing distribution shifts.
@@
Figure 6: Federated data partitioning under different Dirichlet concentration parameter $\alpha$. In each of the subplots, each row represents a client, and each cell corresponds to a data sample whose color indicates its class label. The sub figures are represented as, (a) Sort and partition setting with two fixed classes per client, (b) to (f) Non-IID splits with Dirichlet distribution, having $\alpha\!\to\!\infty$, $\alpha=100,1,0.5$, and $\alpha\!\to\!0$. Larger $\alpha$ produces nearly IID client distributions, whereas smaller $\alpha$ concentrates samples from few classes only within each client with increasingly skewed and heterogeneous partitions.
0.A.2 Federated learning settings
We describe in detail the three personalized Federated Learning: in-distribution (ID), domain generalization (DG) and base-to-new generalization.
[-] In-distribution & Domain generalization. For both settings, we distribute the classes across clients in an overlapping non-independent and identically Distributed (non-IID) manner, as similar to [fedotp]. To simulate controllable non-IID data heterogeneity across clients, we adopt the standard Dirichlet-based partitioning strategy [fedotp, cao2023knowledge, shamsian2021personalized]. Given a dataset with $C$ classes and $N$ clients, each client’s class proportion vector, $\mathbf{q}_{n}=(q_{n1},\dots,q_{nK})$ is sampled from a symmetric Dirichlet distribution [bayesian] parameterized by concentration $\alpha$:
@@ 0.A.2 Federated learning settings
where $|\mathcal{D}_{n}|$ denotes the local dataset size. The concentration parameter $\alpha$ directly controls the heterogeneity of the federated setting i.e. large $\alpha$ values (e.g., $\alpha\!\rightarrow\!\infty$) yield nearly IID client distributions, while small $\alpha$ values (e.g., $\alpha\!\rightarrow\!0$) produce extremely skewed, label-concentrated clients. Figure 6 visualizes this progression, showing how the distribution transitions from highly mixed to strongly partitioned as $\alpha$ decreases. We set $\alpha=0.5$ for both federated settings to maintain a balanced, overlapping non-IID distribution. For in-distribution, we choose 100 clients over a 10% participation rate, whereas two clients per domain with 100% participation rate in domain generalization setting.
@@
0.A.3 Evaluation protocol and metrics
Evaluation protocol. We evaluate the methods in the three aforementioned FL settings in a personalized FL setup [fedotp, fedpha, fedpgp], where we compute the metrics on the test set in each client and then take an average over all the clients. To be noted that client participation in training may vary according to the participation rate, but inference is performed on all clients regardless of their participation.
Metrics. In our FL experiments, we use the accuracy metric for model performance and ECE, MCE, and ACE metrics for model calibration.
@@ 0.A.3 Evaluation protocol and metrics
[-] Expected Calibration Error (ECE). ECE [ece] measures the average discrepancy between a model’s predicted confidence and accuracy. The prediction confidences are partitioned into $G$ static bins, and for each bin $b_{g}$, the absolute difference between the average accuracy $\text{acc}(b_{g})$ and the mean predicted confidence $\text{conf}(b_{g})$ is computed. These bin-wise errors are then weighted by the proportion of samples in each bin, yielding a global estimate of how well the model’s confidence reflects the true likelihood of correctness. Lower ECE values indicate better-calibrated predictions. The ECE metric can be defined as:
@@
[-] Maximum Calibration Error (MCE). MCE [ece] quantifies the worst-case miscalibration across all bins. Rather than averaging, MCE takes the maximum absolute gap between accuracy and confidence over the $G$ static bins. This highlights extreme local over- or under-confidence that may be masked by aggregate metrics such as ECE. A lower MCE value indicates that the model avoids severe calibration failures in any confidence region. The MCE metric can be written as:
@@
[-] Adaptive Calibration Error (ACE). ACE [ace] assesses calibration by computing the discrepancy between predicted confidence and accuracy using adaptive binning. Unlike ECE which divides the predictions into $G$ static bins over fixed confidence intervals, ACE constructs $H$ adaptive bins by sorting the predicted probabilities and dividing them into bins having an equal number of samples, to provide a more realistic calculation of calibration across the entire confidence distribution. ACE evaluates calibration across all class probabilities rather than considering only the maximum predicted confidence. The ACE metric can be defined as:
$\text{ACE}=\frac{1}{C}\frac{1}{H}\sum^{C}_{c=1}\sum^{H}_{h=1}\mid\text{acc}(b_{h,c})-\text{conf}(b_{h,c})\mid,$ (9)
where $\text{acc}(b_{h,c})$ and $\text{conf}(b_{h,c})$ are respectively the average accuracy and predicted probability for class $c$ among the samples belonging to bin $h$.
0.A.4 Calibration baselines
In our experimental settings, we consider two standalone calibration methods: DCA [dca] and MDCA [mdca] as additional plug-ins in the federated methods.
[-] Difference-between-Confidence-Accuracy (DCA). DCA [dca] provides a minibatch-level calibration objective by penalizing the discrepancy between the model’s average predicted confidence score corresponding to the true class and its average accuracy for all the samples in a minibatch. For each prediction corresponding to the $i$-th sample, the method computes the correctness indicator $c_{i}$ (which equals $1$ if the prediction is correct and 0 otherwise) and the predicted confidence of the true label $s_{i}=p(\hat{y}_{i})$. By minimizing this deviation, DCA encourages the model’s overall confidence level to align with its average true accuracy, thereby improving global calibration behavior. For a minibatch of size $m$, the loss is computed as: